OpenStudy (anonymous):

What is the distance between the points (5, –6) and (5, 2)? A. 1 B. 3 C. 4 D. 8

OpenStudy (anonymous):

\[ d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] distance formula

OpenStudy (haseeb96):

\[\sqrt{(x2-x1)^{2}+ (y2-y1)^{2}}\] this is the distance formula now just put these values in it and get the answer (x1,y1)=(5,-6) (x2,y2)=(5,2)

OpenStudy (anonymous):

its 8 option d